OpenStudy (anonymous):

To determine the cost of an in-home repair, a computer technician uses the linear function C(n)=69n+37, where n is the time in hours abd C(n) is the cost in dollars. Fine C(4). For 4 hours the cost is

OpenStudy (anonymous):

C(n) = 69n +37 C(4) : n = 4; C(4) = 69*4 +37 = 313

OpenStudy (anonymous):

plug the 4 in

OpenStudy (anonymous):

oops he beat me to it!